OpenStudy (anonymous):

if m+n is even, then m-n is even use an indirect proof So I assume m-n is odd (indirect proof) m-n=2i+1 where i is Z then m+n = 2i +1 -2(2i+1-m) how do i get the right side to be simplified without ms or ns? help!

OpenStudy (raden):

m+n would be even if m,n both are even or both are odd, now check one by one : even - even = even odd - odd = even QED

OpenStudy (zarkon):

if m-n=2k+1 then m+n=m-n+2n=2k+1+2n=2(k+n)+1 hence m+n is odd (a contradiction)